# Reinforcing Question Answering Agents with Minimalist Policy Gradient Optimization

### v2

@@ 3.1 Preliminary: Reinforcement Learning Setting
If trajectories sampled from $\pi_{\theta}$ are differentiable w.r.t. $\theta$, the gradient of the optimization objective can be estimated using the Vanilla Policy Gradient (VPO) (williams1992simple, ):
@@
3.2 Reinforcement Learning for Large Language Models
Even though the VPO (Eq. 3) estimator is unbiased, it often exhibits high gradient variance, which potentially renders the training process unstable. To address this issue, many previous approaches schulman2015trust ; schulman2017proximal propose to replace the raw trajectory reward $r(\tau)$ with the Generalized Advantage Estimation (GAE) (schulman2016high, ). These RL techniques, particularly Proximal Policy Optimization (PPO) schulman2017proximal and its variants, have been successfully applied in the post-training process for LLM (ouyang2022training, ; rafailov2023direct, ). Recently, Group Relative Policy Optimization (GRPO) (shao2024deepseekmath, ) has been introduced as a simplified alternative to PPO, reducing the training costs while maintaining comparable performance. The objective function of PPO and GRPO, as adapted for our QA context, can be written as follows:
$\mathcal{J}(\theta)=\underset{\tau\sim\pi_{\theta_{k}}}{\mathbb{E}}\left[A(\tau)\sum_{t=0}^{T}\min\left(\mathbf{r}_{t},\;\text{clip}\left(\mathbf{r}_{t},1-\epsilon,1+\epsilon\right)\right)\right]-\beta\mathds{D}_{\text{KL}}(\pi_{\theta}||\pi_{\text{ref}})\;,$ (4)
@@ 3.2 Reinforcement Learning for Large Language Models
3.3 Minimalist Policy Gradient Optimization (MyGO)
Despite its wide adoption, a PPO-style objective has two shortcomings, hyperparameter sensitivity and training inefficiency. The $\epsilon$ hyperparameter is highly sensitive - a small $\epsilon$ may ignore most available trajectories during training, leading to training inefficiency; while a large $\epsilon$ might lose control over the importance sampling scale, resulting in a less stable training process. The advantage estimation renormalizes the original rewards relative to either the value functions or the group average, which weakens the supervision signals from trajectories with large rewards that are helpful for training liu2024decade . It is worth mentioning that training inefficiency and instability are significant challenges when training LLMs, which is both computational expensive and time-consuming. On the other hand, compared with previous RL scenarios in robot control and video games openai2016gym , LLM setting poses new characteristic: fast simulation, deterministic environment, and trajectory-level reward li2024remax . In this scenario, sampling desired trajectories from the environment is relatively more efficient.
To this end, we propose our new RL algorithm, Minimalist Policy Gradient Optimization (MyGO), which shifts the burden from model training to trajectory sampling, in which we can fully leverage the fast and relatively inexpensive simulation of LLMs kwon2023efficient ; zheng2024sglang . Unlike previous methods, MyGO disentangles the RL training process into two phases: sampling and training. During sampling, MyGO directly samples trajectories from the asymptotically optimal distribution and consequently allows the use of Maximum Likelihood Estimation (MLE) to optimize the policy function. Notably, optimizing a model via MLE is simple, stable, and well established as a regular training technique widley used in unsupervised pre-training radford2018improving and supervised fine-tuning ouyang2022training .
@@ 3.3 Minimalist Policy Gradient Optimization (MyGO)
We first characterize the target optimal distribution $\pi^{*}$. Alternative to the original objective $\mathcal{J}$, we consider the expected reward with entropy regularization $\mathcal{J}^{\prime}=\mathbb{E}[r(\tau)]+\alpha\mathbf{H}(\pi)$ with $\alpha\in\mathbb{R}^{+}$, which, as a Boltzmann distribution ziebart2010modeling , has a closed-form for the corresponding optimal policy, denoted as ${\pi^{\prime}}^{*}$:
[…]
@@
In practice, we relax the choice of $K$ by setting an update role of $K:=\max(K^{\prime},(\bar{r}_{n}/r_{\text{sup}}+1)\cdot r_{\text{sup}})$, where $r_{\text{sup}}$ denotes $\sup_{\tau\in\mathcal{T}}r(\tau)$, $\bar{r}_{n}$ is the empirical average reward of a batch of $n$ sampled trajectories, and $K^{\prime}$ is the previous threshold value which is initialized as a moderate value depending on the choice of dataset. As the policy $\pi_{\theta}$ improved, the $K$ will be asymptotically close to $\sup_{\tau\in\mathcal{T}}r(\tau)$, thus making the sampling distribution close to ${\pi^{\prime}}^{*}$.
We highlight that this approach is practically feasible for the reward landscapes for tasks of LLM agent, which differ significantly from that in complex robotic control problems. For LLM agents, trajectories frequently culminate in outcomes that are clearly classifiable as successful or unsuccessful. Empirically, we observe that well-prompted LLMs can readily generate a significant proportion of such successful trajectories, which makes directly sampling from $\pi^{*}$ feasible.
Let $\pi^{>K}$ denote the distribution that consists of trajectories with $r>K$. To clarify, we provide two propositions as follows, where the proofs can be found in the Appendix:
Proposition 3.1.
Given $\delta\in\mathbb{R}$, $\alpha\in\mathbb{R}$, there exists $K\in\mathbb{R}^{+}$ such that the following inequality holds:
$\mathds{D}_{\text{KL}}(\pi^{>K}||\pi^{*})<\delta\;,$
Proposition 3.2.
The following property holds when $\alpha\to 0$:
[…]
Proposition 3.1 states that if $K$ is sufficiently large, that the $\pi^{>K}$ can always approximate $\pi^{*}$. Proposition 3.2 says the variance of $\pi^{>K}$ is small, considering that $Z$ is usually a large value. These two propositions guarantee the proximity and stability of our proposed sampling strategy.
@@
Once a sufficiently large dataset of optimal trajectories $\mathcal{D}$ has been collected, the optimization of policy $\pi_{\theta}$ can be simplified to maximizing the log-likelihood $\mathcal{L}$ of these trajectories. This yields the following MLE gradient:
$\nabla_{\theta}\mathcal{L}$ $=\frac{1}{|\mathcal{D}|}\sum_{\tau_{i}\in\mathcal{D}}\nabla_{\theta}\log\pi_{\theta}(\tau_{i})=\frac{1}{|\mathcal{D}|}\sum_{\tau_{i}\in\mathcal{D}}\sum_{j=0}^{|\tau_{i}|}\nabla_{\theta}\mathds{1}(o_{j}^{i})\log\pi_{\theta}(o_{j}^{i}|o_{<j}^{i})\;,$ (6)
where $o_{j}^{i}$ denotes the $j$-th token from $\tau_{i}$ and $o_{<j}^{i}$ denotes the sequence of tokens preceding $o_{j}$, $\mathds{1}(o_{j})$ is a characteristic function indicating whether $o_{j}$ is a token from the agent. The last equality holds because the action likelihood in language setting could be further elaborated with token likelihood, i.e, $\log\pi_{\theta}(a_{t}|s_{t})=\sum_{j}\log\pi_{\theta}(o_{j}|o_{<j})$.
This MLE objective, Eq. 6, is equivalent to a standard cross-entropy loss function, widely used in both unsupervised pre-training radford2018improving and supervised fine-tuning ouyang2022training of language models. Compared to PPO-style algorithms, Eq. 4, MLE does not incorporate an advantage function, policy ratio clipping, and importance sampling. Similarly, unlike methods such as VPO, Eq. 3, MLE directly uses the curated data $\mathcal{D}$ for training and does not explicitly use reward values in the loss computation. We empirically observed that MLE leads to more stable training. We explain this increased stability by several factors: (i) the dataset $\mathcal{D}$ consists of trajectories sampled from a more stationary (approximated optimal) distribution, reducing variance compared to on-policy RL where the data distribution shifts rapidly; (ii) the direct use of a log-likelihood objective (cross-entropy) is often well-behaved and less sensitive to reward scaling issues that can affect RL algorithms; and (iii) the avoidance of importance sampling ratios, which can introduce instability if they become too large or small. A detailed theoretical justification for these empirical observations is left for future work.
Compared with previous works, like RAFT dongraft and STaR zelikman2022star , who try to simplify RL in LLM post-training, we use a progressive threshold to select data. More importantly, we provide a theoretical justification that elucidates why this approach, yet relatively simple, yields strong performance, offering novel insights into its efficacy.
4 Experiments
4.1 Environment Design
@@
To evaluate the effectiveness of our proposed agent workflow (Mujica) and training method (MyGO) on complex question answering tasks, we utilize four representative benchmark datasets: 2Wiki-MultihopQA (2Wiki) (ho2020constructing, ), QALD-10 (QALD-10, ), HotpotQA (yang2018hotpotqa, ), and MuSiQue (trivedi2021musique, ). 2Wiki and QALD-10 are KBQA benchmarks; HotpotQA and MuSiQue are text-based QA dataset. We use 2Wiki and HotpotQA for training, QALD-10 and MuSiQue for zero-shot evaluation. We assess QA performance using Exact Match (EM) and F1 score. For each dataset, these metrics are computed using the official evaluation scripts provided by the respective dataset creators. Comprehensive dataset statistics and precise definitions of the evaluation metrics are detailed in Appendix.
@@
To evaluate our agent’s capabilities across different data modalities, we derive two experimental environments from the 2Wik dataset: 2Wiki-KG and 2Wiki-Text. This setup allows us to agent’s performance on both structured (KGs) and unstructured (text) data. We use the F1 score as reward function for all environments.
@@
The 2Wiki-KG environment is centered around the Wikidata KG (Wikidata, ). A KG is composed of factual statements represented as triples, such as (Donald Trump, position held, President of the United States). When the agent visits an entity, it can access all triples for which that entity is the head. The agent’s task involves traversing the KG by interpreting the semantic relationships encoded in these triples. For example, to answer the question "Who is the mother of the director of film Polish-Russian War (Film)?", the agent must identify and follow the correct reasoning path, such as: Polish-Russian War, director, Xawery Żuławski, mother, Małgorzata Braunek. To emphasize the reasoning aspect, we follow prior work (ToG, ; ToG2, ) by providing the agent with the gold topic entities extracted from the dataset as the entry.
@@
We establish the 2Wiki-Text environment to further evaluate our agent’s ability to interact with text-based search engines. The original 2Wiki dataset provides ten supporting passages for each question, which contain the necessary information to derive the answer and the "distractor setting" is designed to test multi-hop reasoning abilities without retrieval. In our 2Wiki-Text environment, we adapt this by aggregating all passages to form a large corpus. This approach, following prior work (IRCoT, ; aguerrebere2023similarity, ), simulates an open-domain retrieval task. The agent’s objective is to answer questions by issuing search queries and then synthesizing information from the retrieved passages.
@@
Beyond the 2Wiki environments, we extend our text-based evaluation using the HotpotQA dataset, from which we construct two additional settings: Hotpot and Hotpot-Kimi. In Hotpot, we folow the vanilla setting that uses the Wikipedia dump provided by the dataset yang2018hotpotqa , which only includes the title and the introductory paragraph. However, we found that the retriever karpukhin2020dense is the bottleneck for QA performance in the HotpotQA dataset. To isolate and more directly assess the agent’s reasoning capabilities independent of retrieval imperfections, we introduce the Hotpot-Kimi environment. Similar to but different from the distractor settings (yang2018hotpotqa, ; ho2020constructing, ), in the Hotpot-Kimi environment,the agent is directly provided with the ten gold supporting passages associated with each question, thereby guaranteeing access to all necessary evidence. However, its primary purpose here is to establish an idealized retrieval condition. Therefore, the gold passages are not directly fed into the QA-agent like distractor setting. Detailed examples and prompts for each environment are provided in Appendix.
4.2 Main Results
Table 1: Comparative Performance of Models on 2Wiki-KG and 2Wiki-Text Environments.
@@ 4.2 Main Results
We evaluate our QA agent under three distinct settings: Few-shot, WarmUp, and MyGO. In the Few-shot setting, the agent is provided with a small number of illustrative examples via in-context learning within the prompt. Following common practice, we then perform a WarmUp phase. This phase is to adapt the base model beyond reliance on few-shot exemplars. For this purpose, we fine-tune the model on 1k samples with EM of 1. The MyGO training then commences, using this warmed-up model as its initialization. We experiment with two open-source models: Qwen2.5-7B-Instruct (Qwen) and Llama-3.1-8B-Instruct (Llama).
@@
Table 2: Comparative Performance of Qwen on Hotpot and Hotpot-Kimi Environment.
@@
Table 3 details the results on the HotpotQA dataset. Notably, even the WarmUp models exhibit strong performance, outperforming the few-shot GPT-4.1 baseline. However, the performance gain from MyGO over the WarmUp models is less pronounced in the Hotpot environment. We attribute this observation to two limitations inherent in this setup. Firstly, the efficacy of the dense retriever can be a bottleneck; essential evidence may not be surfaced for some queries. Secondly, the HotpotQA evaluation protocol has no answer aliases. This is problematic because LLMs frequently generate answers that are semantically equivalent to the gold answer but lexically different, which are penalized by strict string-matching metrics. To mitigate the impact of imperfect retrieval, we further evaluate using the Hotpot-Kimi environment, which is designed to simulate near-perfect retrieval by providing gold supporting passages. In this idealized Hotpot-Kimi setting, MyGO demonstrates substantial and consistent performance improvements over the WarmUp models, underscoring its effectiveness when reasoning capabilities can be more directly assessed. A more comprehensive comparison is provided in Appendix.
4.3 Out-Of-Distribution Evaluation
[…]
Type Model H $\to$ M 2W $\to$ Q
EM F1 EM F1
Few-Shot Qwen 2.5-7B 5.34 17.81 27.92 37.11
MyGO 26.11 35.92 39.85 49.73
5 Related Works
MHQA with textual corpus.
Knowledge-based question answering.
6 Conclusion
In this work, we proposed Mujica, an interactive conversational framework for multi-hop question answering, and introduced MyGO, a novel RL optimization method that overcomes the inefficiencies of Proximity Optimization-based approaches. Mujica leverages a planner-worker structure to efficiently decompose and resolve complex questions, while MyGO improves the training efficiency of proximity optimization methods by shifting the burden from clipping importance ratio and tweaking advantage functions to trajectory sampling. Our experiments demonstrate that the proposed framework not only achieves significant performance improvements across multiple QA datasets but also exhibits strong inductive reasoning capabilities, highlighting its scalability and potential to generalize effectively to real-world scenarios.
Appendix A Problem Definition
Text-based Multi-hop Question Answering
Knowledge-based Question Answering
Appendix B Mathematical Proofs
See 3.1
Proof.
[…]
$\mathds{D}_{\text{KL}}(\pi^{>K}||\pi^{*})=\log\frac{Z}{Z^{>K}}\;.$ (8)
Notably,
[…]
See 3.2
Proof.
[…]
Note that we can expand the integral at $\alpha=0$ with a Taylor series:
$\mathrm{Var}[r]=$ (11)
$\frac{e^{(K+M)/\alpha}(-2\alpha^{2}+O(\alpha^{6}))+\left(e^{(2K)/\alpha}+e^{(2M/\alpha)}\right)\left(\alpha^{2}+O(\alpha^{6})\right)+(K-M)^{2}(-e^{(K+M)/\alpha})}{\left(e^{(K/\alpha)}-e^{(M/\alpha)}\right)^{2}}\;.$
Appendix C Datasets and Metrics
To evaluate MyGO’s effectiveness on complex question answering, we employ four representative datasets: 2WikiMultihopQA (2Wiki) [19], QALD-10 [69], HotpotQA [80], and MuSiQue [67]. 2Wiki utilizes both structured (Wikidata) and unstructured (Wikipedia) data to create multi-hop questions. QALD-10 was initially designed for Multilingual Question Answering over KG but late was widely used as a benchmark on KBQA [63]. The other two datasets focus on multi-document and multi-hop reasoning. HotpotQA features complex questions requiring reasoning across multiple documents. MuSiQue consists of questions demanding two to four reasoning hops, constructed from five existing single-hop QA datasets.
Appendix D Implementation Details
D.1 Sampling Strategies
Reward Criterion
Note that some works utilize more complicated non-standard criteria, e.g., LLM-as-Judge [10, 62], Cover Exact Match (CEM) [62, 60]. Since this is not the main focus of this work, we simply follow the standard criterion widely used in previous works.
Progressive Selection Threshold
Multiple-Attempt Sampling
For a batch of $n$ questions, we would iteratively sample trajectories for them. For a question, the sampling procedure will stop if $m$ satisfying trajectories are sampled or reach the predetermined maximum sampling attempts. With vLLM engine [26], we could accelerate the sampling procedures in parallel efficiently with multi-threading/multi-processing.
D.2 Training
Online Off-policy Optimization
In MyGO, we primarily focus on online off-policy optimization. Given a total of $N$ data examples (i.e., questions), we divide them into $T$ batches for $T$ update iterations. For each iteration, we sample trajectories (including answers) from $N/T$ questions to serve as pseudo-labels and optimize the models via MLE through a standard SFT pipeline (e.g., LLaMA-Factory [90]). After completing one update iteration, we proceed to the next sampling iteration.
Trajectory Selection for Training
In the Mujica framework, the planner and workers collaborate to reach the final answer. This interaction generates a single planner trajectory and multiple worker trajectories. It is important to clarify that in Section 3.3, the term "trajectory" specifically refers to the planner’s trajectory. In contrast, the dialogue history for each worker is straightforward, comprising a single-turn interaction with the environment. The environment’s message to the worker consists of the question from the planner and passages from the retriever, while the worker’s message contains its response. If the reward associated with a planner trajectory exceeds a predefined threshold, that trajectory and its corresponding worker trajectories are selected as independent data for training.
Training Advances from Pre-training/SFT
D.3 Hyperparameters and Configurations
Sampling
Warmup
Configuration
Following the settings of regular SFT, we employ the AdamW optimizer [41] with a Cosine Annealing learning rate scheduler [40]. For the warmup phase, we configure the initial learning rate at 1e-5 and the minimum learning rate at 1e-6. In the MyGO training procedure, we set the initial learning rate to 5e-7 and the minimum learning rate to 1e-8. We leverage vLLM [26] as our inference engine and LLaMA-Factory [90] as our training framework.
Appendix E Comprehensive Comparison with Other Methods
Table 4: Evaluation Results on Three Representative Multi-hop Question Answering Datasets.
Method Backbone Retriever HotpotQA MuSiQue 2Wiki
@@ Appendix E Comprehensive Comparison with Other Methods
Concurrent works
@@
Mujica (few-shot) Qwen 2.5-7B bge-large-en-v1.5 15.77 27.96 5.34 17.81 12.41 19.71

### v3

@@ 3.1. Reinforcement Learning Setting
If trajectories sampled from $\pi_{\theta}$ are differentiable w.r.t. $\theta$, the gradient of the optimization objective can be estimated using the Vanilla Policy Gradient (VPO) (williams1992simple):
@@
3.2. Reinforcement Learning for LLMs
Even though the VPO (Eq. 3) estimator is unbiased, it often exhibits high gradient variance, which potentially renders the training process unstable. To address this issue, many previous approaches schulman2015trust; schulman2017proximal propose to replace the raw trajectory reward $r(\tau)$ with the Generalized Advantage Estimation (GAE) (schulman2016high). These RL techniques, particularly Proximal Policy Optimization (PPO) schulman2017proximal and its variants, have been successfully applied in the post-training process for LLM (ouyang2022training; rafailov2023direct). Recently, Group Relative Policy Optimization (GRPO) (shao2024deepseekmath) has been introduced as a simplified alternative to PPO, reducing the training costs while maintaining comparable performance. The objective function of PPO and GRPO, as adapted for our QA context, can be written as follows:
$\mathcal{J}(\theta)=$ $\underset{\tau\sim\pi_{\theta_{k}}}{\mathbb{E}}\left[A(\tau)\sum_{t=0}^{T}\min\left(\mathbf{r}_{t},\;\text{clip}\left(\mathbf{r}_{t},1-\epsilon,1+\epsilon\right)\right)\right]$ (4)
$\hskip 0.0pt-\beta\mathds{D}_{\text{KL}}(\pi_{\theta}||\pi_{\text{ref}})\;,$
@@ 3.2. Reinforcement Learning for LLMs
3.3. MyGO
Despite its wide adoption, a PPO-style objective has two shortcomings: hyperparameter sensitivity and training inefficiency. The $\epsilon$ hyperparameter is highly sensitive - a small $\epsilon$ may ignore most available trajectories during training, leading to training inefficiency; while a large $\epsilon$ might lose control over the importance sampling scale, resulting in a less stable training process. The advantage estimation renormalizes the original rewards relative to either the value functions or the group average, which weakens the supervision signals from trajectories with large rewards that are helpful for training liu2024decade. It is worth mentioning that training inefficiency and instability are significant challenges when training LLMs, which is both computationally expensive and time-consuming. On the other hand, compared with previous RL scenarios in robot control and video games openai2016gym, the LLM setting poses a new characteristic: fast simulation, deterministic environment, and trajectory-level reward li2024remax. In this scenario, sampling desired trajectories from the environment is relatively more efficient.
To this end, we propose our new RL algorithm, MyGO: Minimalist Policy Gradient Optimization, which shifts the burden from model training to trajectory sampling, in which we can fully leverage the fast and relatively inexpensive simulation of LLMs kwon2023efficient; zheng2024sglang. Unlike previous methods, MyGO disentangles the RL training process into two phases: sampling and training. During sampling, MyGO directly samples trajectories from the asymptotically optimal distribution and consequently allows the use of Maximum Likelihood Estimation (MLE) to optimize the policy function. Notably, optimizing a model via MLE is simple, stable, and well established as a regular training technique widely used in unsupervised pre-training radford2018improving and supervised fine-tuning ouyang2022training.
@@ 3.3. MyGO
We first characterize the target optimal distribution $\pi^{*}$. Alternative to the original objective $\mathcal{J}$, we consider the expected reward with entropy regularization $\mathcal{J}^{\prime}=\mathbb{E}[r(\tau)]+\alpha\mathbf{H}(\pi)$ with $\alpha\in\mathbb{R}^{+}$, which, as a Boltzmann distribution ziebart2010modeling, has a closed-form solution for the corresponding optimal policy, denoted as $\pi^{\prime}$$^{*}$:
${\pi^{\prime}}^{*}(\tau)$ $=\frac{\exp(r(\tau)/\alpha)}{Z(\alpha)}\;,$ (5)
$\quad\text{where }Z(\alpha)$ $=\int_{\mathcal{T}}\exp(r(\tau^{\prime})/\alpha)d\tau^{\prime}\;,$
@@
In practice, we relax the choice of $K$ by setting an update role of $K:=\max(K^{\prime},(\bar{r}_{n}/r_{\text{sup}}+1)\cdot r_{\text{sup}})$, where $r_{\text{sup}}$ denotes $\sup_{\tau\in\mathcal{T}}r(\tau)$, $\bar{r}_{n}$ is the empirical average reward of a batch of $n$ sampled trajectories, and $K^{\prime}$ is the previous threshold value which is initialized as a moderate value depending on the choice of dataset. As the policy $\pi_{\theta}$ improved, $K$ will be asymptotically close to $\sup_{\tau\in\mathcal{T}}r(\tau)$, thus making the sampling distribution close to $\pi^{\prime}$$^{*}$.
We highlight that this approach is practically feasible for the reward landscapes for tasks of the LLM agent, which differ significantly from those in complex robotic control problems. For LLM agents, trajectories frequently culminate in outcomes that are clearly classifiable as successful or unsuccessful. Empirically, we observe that well-prompted LLMs can readily generate a significant proportion of such successful trajectories, which makes directly sampling from $\pi^{*}$ feasible.
Let $\pi^{>K}$ denote the distribution that consists of trajectories with $r>K$. To clarify, we provide two propositions as follows, where the proofs can be found in Appendix B: {restatable}propositionklapprox Given $\delta\in\mathbb{R}$, $\alpha\in\mathbb{R}$, there exists $K\in\mathbb{R}^{+}$ such that the following inequality holds:
$\mathds{D}_{\text{KL}}(\pi^{>K}||\pi^{*})<\delta\;$
{restatable}
propositionvarapprox The following property holds when $\alpha$ approaches 0:
$\underset{\pi^{>K}}{\mathrm{Var}[r]}\propto 1/Z^{>K}(\alpha)^{2}\;,\;Z^{>K}(\alpha)=\int^{\sup(r)}_{K}\exp(r/\alpha)dr$
Proposition 3.3 states that if $K$ is sufficiently large, then $\pi^{>K}$ can always approximate $\pi^{*}$. Proposition 3.3 says the variance of $\pi^{>K}$ is small, considering that $Z$ is usually a large value. These two propositions guarantee the proximity and stability of our proposed sampling strategy.
@@
Once a sufficiently large dataset of optimal trajectories $\mathcal{D}$ has been collected, the optimization of policy $\pi_{\theta}$ can be simplified to maximizing the log-likelihood $\mathcal{L}$ of these trajectories. The learning process can be conducted in an online or an offline manner. This yields the following MLE gradient:
$\nabla_{\theta}\mathcal{L}$ $=\frac{1}{|\mathcal{D}|}\sum_{\tau_{i}\in\mathcal{D}}\nabla_{\theta}\log\pi_{\theta}(\tau_{i})$
$=\frac{1}{|\mathcal{D}|}\sum_{\tau_{i}\in\mathcal{D}}\sum_{j=0}^{|\tau_{i}|}\nabla_{\theta}\mathds{1}(o_{j}^{i})\log\pi_{\theta}(o_{j}^{i}|o_{<j}^{i})\;,$ (6)
where $o_{j}^{i}$ denotes the $j$-th token from $\tau_{i}$ and $o_{<j}^{i}$ denotes the sequence of tokens preceding $o_{j}$, $\mathds{1}(o_{j})$ is a characteristic function indicating whether $o_{j}$ is a token from the agent. The last equality holds because the action likelihood in the language setting could be further elaborated with token likelihood, i.e, $\log\pi_{\theta}(a_{t}|s_{t})=\sum_{j}\log\pi_{\theta}(o_{j}|o_{<j})$.
This MLE objective, Eq. 3.3, is equivalent to a standard cross-entropy loss function, widely used in both unsupervised pre-training radford2018improving and supervised fine-tuning ouyang2022training of language models. Compared to PPO-style algorithms, Eq. 4, MLE does not incorporate an advantage function, policy ratio clipping, and importance sampling. Similarly, unlike methods such as VPO, Eq. 3, MLE directly uses the curated data $\mathcal{D}$ for training and does not explicitly use reward values in the loss computation. We empirically observed that MLE leads to more stable training. We explain this increased stability by several factors: (i) the dataset $\mathcal{D}$ consists of trajectories sampled from a more stationary (approximated optimal) distribution, reducing variance compared to on-policy RL where the data distribution shifts rapidly; (ii) the direct use of a log-likelihood objective (cross-entropy) is often well-behaved and less sensitive to reward scaling issues that can affect RL algorithms; and (iii) the avoidance of importance sampling ratios, which can introduce instability if they become too large or small. The implementation details are in Appendix D.
Compared with previous works, like RAFT dongraft and STaR zelikman2022star, which try to simplify RL in LLM post-training, we use a progressive threshold to select data. More importantly, we provide a theoretical justification that elucidates why this approach, yet relatively simple, yields strong performance, offering novel insights into its efficacy.
4. Experiments
4.1. Environment Design
@@
To evaluate the effectiveness of our proposed agent workflow (Mujica) and training method (MyGO) on complex question answering tasks, we utilize four representative benchmark datasets: 2Wiki-MultihopQA (2Wiki) (ho2020constructing), QALD-10 (QALD-10), HotpotQA (yang2018hotpotqa), and MuSiQue (trivedi2021musique). 2Wiki and QALD-10 are KBQA benchmarks; HotpotQA and MuSiQue are text-based QA dataset. We use 2Wiki and HotpotQA for training, QALD-10 and MuSiQue for zero-shot evaluation. We assess QA performance using Exact Match (EM) and F1 score. For each dataset, these metrics are computed using the official evaluation scripts provided by the respective dataset creators. Comprehensive dataset statistics and definitions of the evaluation metrics are detailed in Appendix C. Our experiment is conducted in an online manner. Due to the limited space, we leave the discussion of offline experiment in Appendix E
@@
To evaluate our agent’s capabilities across different data modalities, we derive two experimental environments from the 2Wik dataset: 2Wiki-KG and 2Wiki-Text. This setup allows us to evaluate the agent’s performance on both structured (KGs) and unstructured (text) data. We use the F1 score as the reward function for all environments.
@@
The 2Wiki-KG environment is centered around the Wikidata KG (Wikidata). A knowledge graph (KG) is composed of factual statements represented as triples, such as (Donald Trump, position held, President of the United States). When the agent visits an entity, it can access all triples for which that entity is the head. The agent’s task involves traversing the KG by interpreting the semantic relationships encoded in these triples. For example, to answer the question ”Who is the mother of the director of film Polish-Russian War (Film)?”, the agent must identify and follow the correct reasoning path, such as: Polish-Russian War, director, Xawery Żuławski, mother, Małgorzata Braunek. To emphasize the reasoning aspect, we follow prior work (ToG; ToG2) by providing the agent with the gold topic entities extracted from the dataset as the entry.
@@
We establish the 2Wiki-Text environment to further evaluate our agent’s ability to interact with text-based search engines. The original 2Wiki dataset provides ten supporting passages for each question, which contain the necessary information to derive the answer and the ”distractor setting” is designed to test multi-hop reasoning abilities without retrieval. In our 2Wiki-Text environment, we adapt this by aggregating all passages to form a large corpus. This approach, following prior work (IRCoT; aguerrebere2023similarity), simulates an open-domain retrieval task. The agent’s objective is to answer questions by issuing search queries and then synthesizing information from the retrieved passages.
@@
Beyond the 2Wiki environments, we extend our text-based evaluation using the HotpotQA dataset, from which we construct two additional settings: Hotpot and Hotpot-Kimi. In Hotpot, we follow the vanilla setting that uses the Wikipedia dump provided by the dataset yang2018hotpotqa, which only includes the title and the introductory paragraph. However, we found that the retriever karpukhin2020dense is the bottleneck for QA performance in the HotpotQA dataset. To isolate and more directly assess the agent’s reasoning capabilities independent of retrieval imperfections, we introduce the Hotpot-Kimi environment. Similar to but different from the distractor settings (yang2018hotpotqa; ho2020constructing), in the Hotpot-Kimi environment,the agent is directly provided with the ten gold supporting passages associated with each question, thereby guaranteeing access to all necessary evidence. However, its primary purpose here is to establish an idealized retrieval condition. Therefore, the gold passages are not directly fed into the QA-agent like distractor setting. Detailed examples and prompts for each environment are provided in Appendix G.
4.2. Main Results
4.2.1. Effectieness of MyGO
We evaluate our RAG agent under three distinct settings: Few-Shot, WarmUp, and MyGO. In the Few-Shot setting, the agent is provided with a small number of illustrative examples via in-context learning within the prompt. Following common practice, we then perform a WarmUp phase. This phase is to adapt the base model beyond reliance on few-shot exemplars. For this purpose, we fine-tune the model on 1k samples with EM of 1. The MyGO training then commences, using this warmed-up model as its initialization. We conduct our main experiments with the Qwen2.5-7B-Instruct (Qwen) LLM. The feasibility of using Llama-3.1-8B-Instruct (Llama) as the backbone LLM will also be discussed.
Table 1. Comparative Performance of Models on 2Wiki-KG and 2Wiki-Text Environments.
@@ 4.2.1. Effectieness of MyGO
Table 2. Comparative Performance of Qwen on Hotpot and Hotpot-Kimi Environment.
@@
Table 3. End-to-end textual multi-hop question answering results.
Method Backbone Retriever HotpotQA MuSiQue 2Wiki-Text
@@
RAG-Star Llama 3.1-8B bge-large-en-v1.5 42.0 54.4 13.0 22.2 34.0 42.0
RAG Agents
@@
Mujica (few-shot) Qwen 2.5-7B bge-large-en-v1.5 15.77 27.96 5.34 17.81 23.35 33.52
@@
Table 2 details the results on the HotpotQA dataset. Notably, even the WarmUp models exhibit strong performance, outperforming the few-shot GPT-4.1 baseline. However, the performance gain from MyGO over the WarmUp models is less pronounced in the Hotpot environment. We attribute this observation to two limitations inherent in this setup. Firstly, the efficacy of the dense retriever can be a bottleneck; essential evidence may not be surfaced for some queries. Secondly, the HotpotQA evaluation protocol has no answer aliases. This is problematic because LLMs frequently generate answers that are semantically equivalent to the ground truth but lexically different, which are penalized by strict string-matching metrics. To mitigate the impact of imperfect retrieval, we further evaluate using the Hotpot-Kimi environment, which is designed to simulate near-perfect retrieval by providing gold supporting passages. In this idealized Hotpot-Kimi setting, MyGO demonstrates substantial and consistent performance improvements over the WarmUp models, underscoring its effectiveness when reasoning capabilities can be more directly assessed.
4.2.2. Effectiveness of Mujica
In order to evaluate the contribution of our two critical design choices in the proposed Mujica framework, we replace the planner–worker split with vanilla iterative RAG, and substitute DAG-like reasoning with chain-like reasoning. For the sake of fair comparison, all subsequent experiments are conducted with a pre-trained Qwen 2.5-7B-Instruct model, which eliminates confounding effects introduced by model training.
Table 4. Experimental results for design choices of the proposed Mujica framework.
Settings 2Wiki-KG 2Wiki-Text
EM F1 EM F1
Mujica (Qwen2.5-7B) 57.60 62.24 23.35 33.52
- w/o. planner-worker split 54.85 59.63 23.22 33.40
- w/o. DAG-like reasoning 53.75 59.08 21.73 32.36
From Table 4 we conclude that our proposed DAG-like reasoning paradigm not only enables independent sub-questions to be processed in parallel, but also outperforms the chain-style baseline. We attribute the performance gain to the paradigm’s ability to effectively decompose complex questions and resolve sub-problems through optimal reasoning paths. The structured decomposition allows the model to better capture the underlying logic of multi-hop questions, leading to more accurate and robust reasoning outcomes. In addition, our planner–worker architecture surpasses the vanilla iterative RAG setup. We attribute the better performance to Mujica’s ability to prevent irrelevant information from interfering with the model’s reasoning and planning process.
4.2.3. Mujica-MyGO v.s. Baseline Methods